OpenStudy (zenmo):

Need help with a small simplification. How does 3/2(SqRt 3) = (SqRt 3)/2?

OpenStudy (anonymous):

not clear try the equation tool below

OpenStudy (zenmo):

\[3/2\sqrt{3}=\sqrt{3}/2\]

OpenStudy (anonymous):

multiply top and bottom by \(\sqrt3\)

OpenStudy (anonymous):

you get \[\frac{3\sqrt3}{2\times 3}=\frac{\sqrt3}{2}\]\]

OpenStudy (anonymous):

To rationalize the denominator multiply by \[\frac{ \sqrt{3} }{ \sqrt{3} }\]giving\[\frac{ 3\sqrt{3} }{ 6 }\]\[=\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (zenmo):

Thanks!

OpenStudy (anonymous):

yw